OpenStudy (anitaasks):

How do I calculate the integral from 0 to 3 of (sqrt(9-x^2)+2) using high-school geometry?

OpenStudy (thomas5267):

Are you sure that you can calculate it using high school geometry?

OpenStudy (anitaasks):

Yeah, it is one of the questions on an old midterm

OpenStudy (anitaasks):

Here the answer key's answer, but it makes no sense to me

OpenStudy (perl):

if you set y = sqrt(9-x^2)+2 then y-2 = sqrt(9-x^2) square both sides (y-2)^2 = 9-x^2 x^2 + (y-2)^2 = 9 this is an equation of a circle centered at (0,2) with radius 3

OpenStudy (perl):



OpenStudy (perl):

now the graph in the picture is only this part